OpenStudy (anonymous):

The parent-teacher organization is selling baskets of cookies for a school fundraiser. The materials needed to make each basket cost $3.75, and the baskets are being sold for $10 each. If they spend $75 to advertise their sale, how many baskets must be sold in order to break even?

OpenStudy (anonymous):

75 = (10 - 3.75)*x = 75/6.25 = 12